OpenStudy (anonymous):

what is 26 1/3-5 19/21?

OpenStudy (anonymous):

\[26 \frac{1}{3} - 5 \frac{19}{21}\]I personally hate working with mixed numbers, so I'll rewrite that as \[26 - 5 + \frac{1}{3} - \frac{19}{21}\]Now, rewrite those fractions so that they have the same denominator!

OpenStudy (anonymous):

((((26*3)+1)21)-(((5*21)+19)3))/3*21